OpenStudy (anonymous):

To solve the equation 2sin(2x) = cosx, you should write it as _____. A. cosx(4sinx - 1) = 0 B. 2sinx + cosx = 0 C. cosx(2sinx - 1) = 0

OpenStudy (tkhunny):

\(2\sin(2x) - \cos(x) = 0\) I don't see that option. Use the sine double angle. \(2(2\sin(x)\cos(x)) - \cos(x) = 0\) \(4\sin(x)\cos(x) - \cos(x) = 0\) I don't see that option. Factor. \(\cos(x)(4\sin(x) - 1) = 0\) Now, we're talkin'!!

OpenStudy (anonymous):

How do you solve these kinds of problems?

OpenStudy (tkhunny):

Thought and creativity? Find SOME way to factor.